# Distance-based measures and Epsilon-measures for measurement-based quantum resources

### v3

@@ III.1 Transformation of sets of measurements, sets of channels and a distance measure
where $[\overline{\mathcal{V}}(\mathcal{C})]_{j}$ is the $j$th channel of the set $\overline{\mathcal{V}}(\mathcal{C})$, $\Sigma_{\mathcal{C}}=\sum_{\overline{i}}\Lambda_{\overline{i}}\otimes\overline{\Phi}_{\overline{i}}$, $|\overline{\Phi}_{\overline{i}}(.)=\ket{\overline{i}}\bra{\overline{i}}(.)\ket{\overline{i}}\bra{\overline{i}}$ | for all $\overline{i}$, $\overline{\Theta}_{pre}^{j}\in\mathscr{C}({\overline{\mathcal{H}},\mathcal{H}\otimes\mathcal{H}_{I}\otimes\overline{\mathcal{R}}})$ for all $j$, $\overline{\Theta}_{post}^{j}\in\mathscr{C}({\mathcal{K}\otimes\mathcal{H}_{I}\otimes\overline{\mathcal{R}},\overline{\mathcal{K}}})$ for all $j$, {|$\{\ket{\overline{i}}\}$ is an orthonormal basis of Hilbert space $\mathcal{H}_{I}$. Note that $\Sigma_{\mathcal{C}}$ is the controlled implementation of the channels in the set $\mathcal{C}$. Clearly, if both $\mathcal{C}$ and $\mathcal{V}[\mathcal{C}]$ contain one quantum channel then (26) reduces to (3) as a special case.
@@
For two sets of measurements $\overline{\mathcal{D}}(\mathcal{C}_{1},\mathcal{C}_{2})$ is a valid distance.
@@
Thirdly, consider another set of channels $\mathcal{C}_{3}=\{\Psi_{i}\}$ now, suppose that in Eq. (27) supremum occurs for $i=i^{*}$ then
@@
The tensor product of two sets of channels $\tilde{\mathcal{C}}_{1}=\{\tilde{\Lambda}_{i}\in\mathscr{C}(\mathcal{H},\mathcal{K}_{1})\}^{m}_{i=1}$ and $\tilde{\mathcal{C}}_{2}=\{\tilde{\Theta}_{j}\in\mathscr{C}(\mathcal{H},\mathcal{K}_{2})\}^{n}_{j=1}$ is defined as $\tilde{\mathcal{C}}_{1}\otimes\tilde{\mathcal{C}}_{2}:=\{\tilde{\Psi}_{ij}=\tilde{\Lambda}_{i}\otimes\tilde{\Theta}_{j}\}$. The convex combination of two sets of channels $\mathcal{C}_{1}=\{\Lambda_{i}\in\mathscr{C}(\mathcal{H},\mathcal{K})\}^{n}_{i=1}$ and $\mathcal{C}_{2}=\{\Theta_{i}\in\mathscr{C}(\mathcal{H},\mathcal{K})\}^{n}_{i=1}$ is defined as $p\mathcal{C}_{1}+(1-p)\mathcal{C}_{2}:=\{\Psi_{i}=p\Lambda_{i}+(1-p)\Theta_{i}\}$. We denote a process where a given set of channels $\mathcal{C}_{1}$ (as an input) is probabilistically mixed with an arbitrary fixed set of channels $\mathcal{Q}=\{\Lambda^{\prime}_{i}\}$ as
@@
$\overline{\mathcal{D}}(\overline{\mathcal{V}}(\mathcal{C}_{1}),\overline{\mathcal{V}}(\mathcal{C}_{2}))=\mathcal{D}_{\Diamond}([\overline{\mathcal{V}}(\mathcal{C}_{1})]_{j^{*}},[\overline{\mathcal{V}}(\mathcal{C}_{2}))]_{j^{*}}).$ (38)
@@
$\overline{\mathcal{D}}(\overline{\mathcal{V}}(\mathcal{C}_{1}),\overline{\mathcal{V}}(\mathcal{C}_{2}))$
$=$ $\mathcal{D}_{\Diamond}([\overline{\mathcal{V}}(\mathcal{C}_{1})]_{j^{*}},[\overline{\mathcal{V}}(\mathcal{C}_{2}))]_{j^{*}})$
$=$ $\mid\mid[\overline{\mathcal{V}}(\mathcal{C}_{1})]_{j^{*}}-[\overline{\mathcal{V}}(\mathcal{C}_{2}))]_{j^{*}}\mid\mid_{\Diamond}$
@@
$=$ $\max_{\overline{\rho}_{\mathcal{H}\mathcal{H}_{I}\mathcal{R}}\in\mathcal{S}(\mathcal{H}\otimes\mathcal{H}_{I}\otimes\mathcal{R})}\mid\mid(\Sigma_{\mathcal{C}_{1}}-\Sigma_{\mathcal{C}_{2}})\otimes\mathbbm{I_{\mathcal{R}}}((\overline{\rho}_{\mathcal{H}\mathcal{H}_{I}\mathcal{R}}))\mid\mid_{1}$ (39)
@@
$=$ $\sum_{i}\mid\mid(\Lambda_{i}-\Theta_{i})\otimes\mathbbm{I}_{\mathcal{R}}(\bra{i}\rho_{\mathcal{H}\mathcal{H}_{I}\mathcal{R}}\ket{i})\mid\mid_{1}$
@@
Now, we define the distance between two sets of measurements =$\mathcal{M}=\{M_{i}\}$ and $\mathcal{N}=\{N_{i}\}$ as
@@
It should be mentioned here that in Ref. [23], authors studied the distance measure for a set of measurements and they defined the distance between two sets of measurements =$\mathcal{M}=\{M_{i}\}$ and $\mathcal{N}=\{N_{i}\}$ as $\mathbf{D}_{\Diamond}(\mathcal{M},\mathcal{N})=\sum_{i}p_{i}\mathcal{D}_{\Diamond}(\Gamma_{M_{i}},\Gamma_{N_{i}})$ where $\{p_{i}\}$ is an arbitrarily chosen probability distribution. Instead of using that measure, in this work, we use the distance measure $\widetilde{\mathcal{D}}$ which is independent of any such probability distribution and is a function of only the sets $\mathcal{M}$ and $\mathcal{N}$. Furthermore, neither $\mathbf{D}_{\Diamond}$ in Ref. [23] nor $\widetilde{\mathcal{D}}$ is independent of permutation of the measurements in the sets. However, that does not affect any of our results and the resource measure created from $\mathbbm{D}$ (which is more general than $\widetilde{\mathcal{D}}$) is independent of the permutation of the measurements as observed from Eq. (55).
@@
Now, consider two sets of measurements =$\mathcal{M}=\{M_{i}\in\mathscr{M}(\mathcal{H})\}$ and $\mathcal{N}=\{N_{i}\in\mathscr{M}(\mathcal{H})\}$ where $\Omega_{M_{i}}=\Omega_{N_{i}}=X_{i}\times Y_{i}\leavevmode\nobreak\ \forall i$ where $X_{i}$s and $Y_{i}$s are some sets of natural numbers. Then define the action of the map $\mathbbm{Merg}_{\mathcal{Y}}$ on $\mathcal{M}$ as
@@
$\tilde{\mathcal{D}}$ is non-increasing under merginalization. In other words, for two sets of measurements $\mathcal{M}$ and $\mathcal{N}$
@@
  • Consider a set of measurements =$\mathcal{M}=\{M_{i}\in\mathscr{M}(\mathcal{H})\}$. An arbitrary free operation $\mathcal{W}_{incomp}$ of the resource theory of measurement incompatibility acting on $\mathcal{M}$ can be written as
@@
    where $\mathcal{I}_{\mu}=\{\tilde{\Phi}_{l|\mu}:\mathcal{L}(\mathcal{K})\rightarrow\mathcal{L}(\mathcal{H})\}$ are the quantum instruments for all $\mu$, $p(\mu|j)$, $r(y|i,x,\mu,j,l)$, and $q(i|\mu,j,l)$ with $p(\mu|j)$, $r(y|i,x,\mu,j,l)$, and $q(i|\mu,j,l)$ being the conditional probabilities [12]. Clearly, $\mathcal{W}_{incomp}[\mathcal{M}]_{j}$ is the $j$-th measurement of the transformed set $\mathcal{W}_{incomp}[\mathcal{M}]\subset\mathscr{M}(\mathcal{K})$. Now, consider the quantum instrument $\mathcal{J}=\{\Phi_{\lambda}=p(\mu)\tilde{\Phi}^{\dagger}_{l|\mu}\}$ where $\lambda=(\mu,l)$ Then the eq. (46) can be simplified as
@@
    where $\mathcal{E}(\rho)\in\mathscr{C}(\mathcal{K},\mathcal{H}\otimes\mathcal{H}_{\Lambda})$ such that $\mathcal{E}(\rho)=\sum_{\lambda}\Phi_{\lambda}(\rho)\otimes\ket{\lambda}\bra{\lambda}$ | for all $\rho\in\mathcal{L}(\mathcal{K})$ where {|$\{\ket{\lambda}\}$ forms an orthonormal basis of Hilbert space $\mathcal{H}_{\Lambda}$, for all $j$, $\Psi_{j}\in\mathscr{C}(\mathcal{H}\otimes\mathcal{H}_{\Lambda},\mathcal{H}\otimes\mathcal{H}_{I}\otimes\mathcal{H}_{\Lambda})$ such that $\Psi_{j}(\sigma)=\sum_{i,\lambda}q(i\mid\lambda,j)\ket{i,\lambda}\bra{\lambda}\sigma\ket{\lambda}\bra{i,\lambda}$ | for all $\rho\in\mathcal{L}(\mathcal{H}\otimes\mathcal{H}_{\Lambda})$ where {|$\{\ket{i}\}$ forms an orthonormal basis of Hilbert space $\mathcal{H}_{I}$, and for all $j$, $\Theta_{j}\in\mathscr{C}(\mathcal{H}_{A}\otimes\mathcal{H}_{I}\otimes\mathcal{H}_{\Lambda},\mathcal{H}_{B})$ such that $\Theta_{j}(\omega)=\sum_{y,x,i,\lambda}r(y|i,x,\lambda,j)\ket{y}\bra{x,i,\lambda}\omega\ket{x,i,\lambda}\bra{y}$ | for all $\omega\in\mathcal{L}(\mathcal{H}_{A}\otimes\mathcal{H}_{I}\otimes\mathcal{H}_{\Lambda})$ where {|$\{\ket{y}\}$ forms an orthonormal basis of Hilbert space $\mathcal{H}_{B}$.
@@
  • Consider a measurement $M=\{M(x)\}\in\mathscr{M}(\mathcal{H})$. An arbitrary free operation $\mathcal{W}_{sharp}$ of the resource theory of measurement sharpness acting on $M$ can be written as
    $\mathcal{W}_{sharp}[M](x)=\mu\Lambda^{\dagger}(M(x))+(1-\mu)p(x)\mathbbm{1}_{\mathcal{H}},$ (50)
@@
    Now, let $N:=\{N(x)=p(x)\mathbbm{1}_{\mathcal{H}}\}$. Then
@@
  • An arbitrary free operation $\mathcal{W}_{coh}$ of the resource theory of measurement coherence acting on $M$ can be written as
@@
    where $K_{j}=\sum_{k}c_{k,i}\ket{\pi_{k}(i)}\bra{i}$ |, $\pi_{k}$ is the permutation, {|$\{\ket{i}\}$ is the orthonormal basis, and $\sum_{j}K^{\dagger}_{j}K_{j}=\mathbbm{1}_{\mathcal{H}}$ [15]. Clearly, $\Gamma_{\mathcal{W}_{coh}[M]}=\Gamma_{M}\circ\Lambda_{coh}$. The pre-processing $\Lambda_{coh}$ is a special case of eq. (26). Therefore, from Theorem 1, and Eq. (41), we obtain
@@
Therefore, $\widetilde{\mathcal{D}}$ is monotononically non-increasing under free operations of (1) the resource theory of measurement incompatibility, (2) the resource theory of measurement sharpness, and (3) the resource theory of measurement coherence.
@@
    $\mathbbm{D}$ is non-increasing (at least) under the free operation of a given resource theory.
@@ III.2 Distance-based measures for measurement-based resources
Consider a set of measurements =$\mathcal{M}=\{M_{i}\in\mathscr{M}(\mathcal{H}_{A})\}$. Then the enlarged version of $\mathcal{M}$ is defined as ^$\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}=\{(\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\}$. Clearly, ^$\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}=\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}}$. Here, $\mathcal{T}_{\mathcal{H}_{B}}$ is considered as a set with one element.
@@
¯$\overline{\mathbbm{R}}(\mathcal{M})=\inf_{\mathcal{H}_{B}}\min_{\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\in\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}\mathbbm{D}(\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}),$ (55)
where $\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}$ is the set of free sets of measurements acting on the Hilbert space $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ and $\mathcal{H}_{B}$ with $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ and $\mathcal{H}_{B}$ being an arbitrary finite dimensional Hilbert space. We have used the suffix $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ to denote the Hilbert space the measurements act on. Throughout this subsection, we will stick to this particular notation.
@@
where $\mathcal{N}^{*}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\in\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}$ is a free set of measurements. Note that ¯$\overline{\mathbbm{R}}(\mathcal{M})\geq 0$ for all $\mathcal{M}$ as $\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})\geq 0$ for all $\mathcal{M}$ and for all $\mathcal{H}_{B}$. Now, if $\mathcal{M}\notin\mathcal{F}_{\mathcal{H}_{A}}$, then by assumption [A1], we have ^$\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\notin\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\leavevmode\nobreak\ \forall\mathcal{H}_{B}$. Then for all $\mathcal{H}_{B}$, we have $\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})>0$ and therefore, from Eq. 55, we have ¯$\overline{\mathbbm{R}}>0$. Now, assume, $\mathcal{M}\in\mathcal{F}_{\mathcal{H}_{A}}$. Then as $\mathbbm{D}(\mathcal{M},\mathcal{M})=0$ from Eq. 55, we have ¯$\overline{\mathbbm{R}}=0$.
Now, for an arbitrary $\mathcal{H}_{B}$
@@
$\geq$ $\min_{\mathcal{N}_{\mathcal{H}_{B}}\in\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}\mathbbm{D}(\hat{\mathcal{W}[\mathcal{M}]}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
@@
where in second line, we have used the fact that $\mathcal{W}\otimes\mathcal{I}^{\dagger}_{\mathcal{H}_{B}}$ is a free transformation as $\mathcal{W}$ is a free transformation as given by assumptions [A2] and [A3], and in the third line, we used the obvious fact that $\mathcal{W}\otimes\mathcal{I}^{\dagger}_{\mathcal{H}_{B}}[\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}]=\hat{\mathcal{W}[\mathcal{M}]}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}$ as ^$\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}=\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}}$. Therefore,
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})\geq$ $\mathbbm{K}(\mathcal{W}[\mathcal{M}],\mathcal{H}_{B})\leavevmode\nobreak\ \forall\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ \inf_{\mathcal{H}_{B}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})\geq$ $\inf_{\mathcal{H}_{B}}\mathbbm{K}(\mathcal{W}[\mathcal{M}],\mathcal{H}_{B})$
$or,\leavevmode\nobreak\ \overline{\mathbbm{R}}(\mathcal{M})\geq$$\overline{\mathbbm{R}}(\mathcal{W}[\mathcal{M}]).$ (60)
@@
¯$\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}})=\overline{\mathbbm{R}}(\mathcal{M})$ for any =$\mathcal{M}=\{M_{i}\in\mathscr{M}(\mathcal{H}_{A})\}$.

### v5

@@ III.1 Transformation of sets of measurements, sets of channels and a distance measure
[…]
Figure 1: Schematic circuit diagram of the fairly general transformation of a set of channels $\mathcal{C}$ to another set of channels $\overline{\mathcal{V}}(\mathcal{C})$ as represented through Eq. (26). In this figure, instead of the actual states of the input and output systems, we mostly indicate input and output Hilbert spaces because it is more important for our purpose of illustration.
@@
For two sets of channels $\overline{\mathcal{D}}(\mathcal{C}_{1},\mathcal{C}_{2})$ is a valid distance.
@@
Thirdly, consider another set of channels $\mathcal{C}_{3}=\{\Psi_{i}\}$. Now, suppose that in Eq. (27) supremum occurs for $i=i^{*}$ then
@@
The tensor product of two sets of channels $\tilde{\mathcal{C}}_{1}=\{\tilde{\Lambda}_{i}\in\mathscr{C}(\mathcal{H}_{1},\mathcal{K}_{1})\}^{m}_{i=1}$ and $\tilde{\mathcal{C}}_{2}=\{\tilde{\Theta}_{j}\in\mathscr{C}(\mathcal{H}_{2},\mathcal{K}_{2})\}^{n}_{j=1}$ is defined as $\tilde{\mathcal{C}}_{1}\otimes\tilde{\mathcal{C}}_{2}:=\{\tilde{\Psi}_{ij}=\tilde{\Lambda}_{i}\otimes\tilde{\Theta}_{j}\}$. The convex combination of two sets of channels $\mathcal{C}_{1}=\{\Lambda_{i}\in\mathscr{C}(\mathcal{H},\mathcal{K})\}^{n}_{i=1}$ and $\mathcal{C}_{2}=\{\Theta_{i}\in\mathscr{C}(\mathcal{H},\mathcal{K})\}^{n}_{i=1}$ is defined as $p\mathcal{C}_{1}+(1-p)\mathcal{C}_{2}:=\{\Psi_{i}=p\Lambda_{i}+(1-p)\Theta_{i}\}$. We denote a process where a given set of channels $\mathcal{C}_{1}$ (as an input) is probabilistically mixed with an arbitrary fixed set of channels $\mathcal{Q}=\{\Lambda^{\prime}_{i}\}$ as
@@
$\overline{\mathcal{D}}(\overline{\mathcal{V}}(\mathcal{C}_{1}),\overline{\mathcal{V}}(\mathcal{C}_{2}))=\mathcal{D}_{\Diamond}([\overline{\mathcal{V}}(\mathcal{C}_{1})]_{j^{*}},[\overline{\mathcal{V}}(\mathcal{C}_{2})]_{j^{*}}).$ (38)
@@
$\overline{\mathcal{D}}(\overline{\mathcal{V}}(\mathcal{C}_{1}),\overline{\mathcal{V}}(\mathcal{C}_{2})$
$=$ $\mathcal{D}_{\Diamond}([\overline{\mathcal{V}}(\mathcal{C}_{1})]_{j^{*}},[\overline{\mathcal{V}}(\mathcal{C}_{2})]_{j^{*}})$
$=$ $\mid\mid[\overline{\mathcal{V}}(\mathcal{C}_{1})]_{j^{*}}-[\overline{\mathcal{V}}(\mathcal{C}_{2})]_{j^{*}}\mid\mid_{\Diamond}$
@@
$=$ $\max_{\overline{\rho}_{\mathcal{H}\mathcal{H}_{I}\mathcal{R}}\in\mathcal{S}(\mathcal{H}\otimes\mathcal{H}_{I}\otimes\mathcal{R})}\mid\mid(\Sigma_{\mathcal{C}_{1}}-\Sigma_{\mathcal{C}_{2}})\otimes\mathbbm{I_{\mathcal{R}}}((\overline{\rho}_{\mathcal{H}\mathcal{H}_{I}\mathcal{R}})\mid\mid_{1}$ (39)
@@
[…]
@@
Now, we define the distance between two sets of measurements $\mathcal{M}=\{M_{i}\}$ and $\mathcal{N}=\{N_{i}\}$ as
@@
It should be mentioned here that in Ref. [23], authors studied the distance measure for a set of measurements and they defined the distance between two sets of measurements $\mathcal{M}=\{M_{i}\}$ and $\mathcal{N}=\{N_{i}\}$ as $\mathbf{D}_{\Diamond}(\mathcal{M},\mathcal{N})=\sum_{i}p_{i}\mathcal{D}_{\Diamond}(\Gamma_{M_{i}},\Gamma_{N_{i}})$, where $\{p_{i}\}$ is an arbitrarily chosen probability distribution. Furthermore, the monotonicity of the resource measure (constructed from their distance measure) under the free transformation of the resource theory of incompatibility of measurements (given in Ref. [12]) has not been shown yet. Instead of using that measure, in this work, we use the distance measure $\widetilde{\mathcal{D}}$ which is independent of any such probability distribution and is a function of only the sets $\mathcal{M}$ and $\mathcal{N}$. Furthermore, neither $\mathbf{D}_{\Diamond}$ in Ref. [23] nor $\widetilde{\mathcal{D}}$ is independent of permutation of the measurements in the sets. However, that does not affect any of our results and the resource measure created from $\mathbbm{D}$ (which is more general than $\widetilde{\mathcal{D}}$) is independent of the permutation of the measurements as observed from Eq. (55).
@@
Now, consider two sets of measurements $\mathcal{M}=\{M_{i}\in\mathscr{M}(\mathcal{H})\}$ and $\mathcal{N}=\{N_{i}\in\mathscr{M}(\mathcal{H})\}$ where $\Omega_{M_{i}}=\Omega_{N_{i}}=X_{i}\times Y_{i}\penalty 10000\ \forall i$ where $X_{i}$s and $Y_{i}$s are some sets of natural numbers. Then define the action of the map $\mathbbm{Merg}_{\mathcal{Y}}$ on $\mathcal{M}$ as
@@
$\widetilde{\mathcal{D}}$ is non-increasing under merginalization. In other words, for two sets of measurements $\mathcal{M}$ and $\mathcal{N}$
@@
  • Consider a set of measurements $\mathcal{M}=\{M_{i}\in\mathscr{M}(\mathcal{H})\}$. An arbitrary free transformation $\mathcal{W}_{incomp}$ of the resource theory of measurement incompatibility acting on $\mathcal{M}$ can be written as
@@
    where $\mathcal{I}_{\mu}=\{\tilde{\Phi}_{l|\mu}:\mathcal{L}(\mathcal{K})\rightarrow\mathcal{L}(\mathcal{H})\}$ are the quantum instruments for all $\mu$, with $p(\mu|j)$, $r(y|i,x,\mu,j,l)$, and $q(i|\mu,j,l)$ being the conditional probabilities [12]. Clearly, $\mathcal{W}_{incomp}[\mathcal{M}]_{j}$ is the $j$-th measurement of the transformed set $\mathcal{W}_{incomp}[\mathcal{M}]\subset\mathscr{M}(\mathcal{K})$. Now, consider the quantum instrument $\mathcal{J}=\{\Phi_{\lambda}=p(\mu)\tilde{\Phi}^{\dagger}_{l|\mu}\}$ where $\lambda=(\mu,l)$ Then the eq. (46) can be simplified as
@@
[…]
@@
  • Consider a measurement $M=\{M(x)\}\in\mathscr{M}(\mathcal{H})$. An arbitrary free transformation $\mathcal{W}_{sharp}$ of the resource theory of measurement sharpness acting on $M$ can be written as
    $\mathcal{W}_{sharp}[M](x)=\mu\Lambda^{\dagger}(M(x))+(1-\mu)p(x)\mathbbm{1}_{\mathcal{K}},$ (50)
@@
    Now, let $N:=\{N(x)=p(x)\mathbbm{1}_{\mathcal{K}}\}$. Then
@@
  • An arbitrary free transformation $\mathcal{W}_{coh}$ of the resource theory of measurement coherence acting on $M$ can be written as
@@
[…]
@@
Therefore, $\widetilde{\mathcal{D}}$ is monotononically non-increasing under free transformation of (1) the resource theory of measurement incompatibility, (2) the resource theory of measurement sharpness, and (3) the resource theory of measurement coherence.
@@
    $\mathbbm{D}$ is non-increasing (at least) under the free transformation of a given resource theory.
@@ III.2 Distance-based measures for measurement-based resources
Consider an arbitrary set of measurements $\mathcal{M}=\{M_{i}\in\mathscr{M}(\mathcal{H}_{A})\}$. Then the enlarged version of $\mathcal{M}$ is defined as $\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}=\{(\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\}$. Clearly, $\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}=\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}}$. Here, $\mathcal{T}_{\mathcal{H}_{B}}$ is considered as a set with one element.
@@
$\overline{\mathbbm{R}}(\mathcal{M})=\inf_{\mathcal{H}_{B}}\min_{\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\in\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}\mathbbm{D}(\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}),$ (55)
where $\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}$ is the set of free sets of measurements (obviously with the same number of elements as the set $\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}$) acting on the Hilbert space $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ and $\mathcal{H}_{B}$ being an arbitrary finite dimensional Hilbert space. We have used the suffix $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ to denote the Hilbert space the measurements act on. Throughout this subsection, we will stick to this particular notation.
@@
where $\mathcal{N}^{*}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\in\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}$ is a free set of measurements. Note that $\overline{\mathbbm{R}}(\mathcal{M})\geq 0$ for all $\mathcal{M}$ as $\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})\geq 0$ for all $\mathcal{M}$ and for all $\mathcal{H}_{B}$. Now, if $\mathcal{M}\notin\mathcal{F}_{\mathcal{H}_{A}}$, then by assumption [A1], we have $\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\notin\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\penalty 10000\ \forall\mathcal{H}_{B}$. Then for all $\mathcal{H}_{B}$, we have $\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})>0$ and therefore, from Eq. 55, we have $\overline{\mathbbm{R}}>0$. Now, assume, $\mathcal{M}\in\mathcal{F}_{\mathcal{H}_{A}}$. Then as $\mathbbm{D}(\mathcal{M},\mathcal{M})=0$ from Eq. 55, we have $\overline{\mathbbm{R}}=0$.
Consider a free transformation $\mathcal{W}$ that transforms the set of measurements $\mathcal{M}\subset\mathscr{M}(\mathcal{H}_{A})$ to another set of measurements $\mathcal{W}[M]\subset\mathscr{M}(\mathcal{K}_{A})$. Now, for an arbitrary $\mathcal{H}_{B}$
@@
$\geq$ $\min_{\mathcal{N}_{\mathcal{H}_{B}}\in\mathcal{F}_{\mathcal{K}_{A}\otimes\mathcal{H}_{B}}}\mathbbm{D}(\hat{\mathcal{W}[\mathcal{M}]}_{\mathcal{K}_{A}\otimes\mathcal{H}_{B}},\mathcal{N}_{\mathcal{K}_{A}\otimes\mathcal{H}_{B}})$
@@
where in second line, we have used the fact that $\mathcal{W}\otimes\mathcal{I}^{\dagger}_{\mathcal{H}_{B}}$ is a free transformation as $\mathcal{W}$ is a free transformation as given by assumptions [A2] and [A3], and in the third line, we used the obvious fact that $\mathcal{W}\otimes\mathcal{I}^{\dagger}_{\mathcal{H}_{B}}[\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}]=\hat{\mathcal{W}[\mathcal{M}]}_{\mathcal{K}_{A}\otimes\mathcal{H}_{B}}$ as $\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}=\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}}$. Therefore,
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})\geq$ $\mathbbm{K}(\mathcal{W}[\mathcal{M}],\mathcal{H}_{B})\penalty 10000\ \forall\mathcal{H}_{B}$
$or,\penalty 10000\ \inf_{\mathcal{H}_{B}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})\geq$ $\inf_{\mathcal{H}_{B}}\mathbbm{K}(\mathcal{W}[\mathcal{M}],\mathcal{H}_{B})$
$or,\penalty 10000\ \overline{\mathbbm{R}}(\mathcal{M})\geq$ $\overline{\mathbbm{R}}(\mathcal{W}[\mathcal{M}]).$ (60)
@@
$\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}})=\overline{\mathbbm{R}}(\mathcal{M})$ for any $\mathcal{M}=\{M_{i}\in\mathscr{M}(\mathcal{H}_{A})\}$.